2\3(7\8-4x)-5\8=3/8
Answer by algebrahouse.com(1659)

"2\3(7\8 - 4x) - 5\8 = 3/8"

14/24 - 8x/3 - 5/8 = 3/8 {used distributive property with 2/3}
24[14/24 - 8x/3 - 5/8 = 3/8] {multiplied entire equation by 24 to clear fractions}
24(14/24) - 24(8x/3) - 24(5/8) = 24(3/8){multiplied each term by 24}
14 - 64x - 15 = 9 {cancelled and multiplied what's left}
-64x - 1 = 9 {combined like terms}
-64x = 10 {added 1 to both sides}
x = -10/64 {divided both sides by -64}
x = -5/32 {reduced}

www.algebrahouse.com